# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ III.4 Electronic structures: band structure
Figure 4: Calculated unfolded band dispersion for $2\times 2$ ISD phase along K-M-K path for different values of $U$ and $V$ (in eV): (a) cRPA values ($U=1.35,V=0.17$), (b) $U=0$, $V=0$, (c) $U=1$, $V=0$, (d) $U=1.5$, $V=0$, (f) $U=1.5$, $V=0.5$, (g) $U=1.5$, $V=1$, and (h) $U=1.5$, $V=1.5$. (e) ARPES intensity measured at $T$ = 20 K along the K-M-K cut at $h\nu=114$ eV, as adapted from Kato et al. Kato et al. (2022). The black solid lines in (a) are a guide for the eye to understand both the SP1 and SP2 bands. A Fermi level offset of 200 meV is introduced in the DFT+$U$+$V$ bands to enable direct comparison with the experimental spectra.
We now investigate the electronic structure of KV3S5, focusing on its dependence on the Hubbard parameters $U$ and $V$. Given the weak interlayer interactions, we consider the $2\times 2$ ISD phase, as shown in Fig. 1. To isolate the effects of $U$ and $V$ from structural contributions, we use the optimized ISD structure obtained at $U=0$ and $V=0$ as the baseline for calculations with various $U$ and $V$ values. The phase shifts between adjacent kagome layers have minimal impact on the band dispersion. Thus, our approach is sufficient for investigating the electronic structure of KV3S5. Structural contributions, such as variations in vanadium atom distortion for different values of $U$ and $V$, will be discussed at the end of this subsection.
Figure 4 presents the band dispersion of the $2\times 2$ ISD phase for various values of $U$ and $V$, unfolded onto the $1\times 1$ Brillouin zone. The $k$-path K-M-K is selected to facilitate direct comparison with existing angle-resolved photoemission spectroscopy (ARPES) measurements, as shown in Fig. 4(e).
The band dispersion of the ISD phase successfully captures two dominant features near the Fermi level: (i) an M-shaped band and (ii) a $\Lambda$-shaped band, as highlighted by the guidelines in Fig. 4(a). These features are consistent with the experimental ARPES measurement by Kato et al. Kato et al. (2022), where the former is denoted as the SP1 band and the latter as SP2. Near the M-point, the $\Lambda$-shaped (SP2) band penetrates the M-shaped (SP1) band and reaches a higher binding energy in the ISD phase. In contrast, no such penetration occurs in the SD phase, where the $\Lambda$-shaped (SP2) band remains at a higher energy position.
We found that our on-site and inter-site Hubbard parameters, $U$ = 1.35 eV and $V$ = 0.17 eV, obtained from the cRPA calculations, successfully reproduce the key features of the experimental ARPES described above (Fig. 4 (a)). Both the relative energy positions of the SP1 and SP2 bands and their morphology closely match the experimental spectra in Fig. 4(e). This agreement contrasts with the results obtained using other $U$ and $V$ values, as shown in the Fig. 4. While DFT calculations without $U$ and $V$ (Fig. 4(b)) also show reasonable agreement with ARPES data, the cRPA-derived finite $U$ and $V$ values play a crucial role in shaping the band dispersion of the KV3Sb5 system, as demonstrated in Fig. S5 of the Supplementary Material Sup .
Our systematic investigation reveals the distinct roles of $U$ and $V$ in shaping the low-energy band structures. Variations in $U$ primarily affect the M-shaped (SP1) band, which predominantly has $3d_{x^{2}-y^{2}}$, $3d_{xy}$, and $3d_{z^{2}}$ orbital characters,
while the $\Lambda$-shaped (SP2) band, mainly associated with $3d_{xz/yz}$ orbital characters, remains nearly rigid (see Fig. S7 in the Supplementary Material Sup for detailed orbital characters). Note that the orbital characters of both SP1 and SP2 bands are consistent with the earlier findings by Kato et al. Kato et al. (2022). Increasing $U$ progressively shifts the SP1 band downward, lowering its energy at the $K$-point, whereas $U$ has minimal impact on the position of the SP2 band. In contrast, the effect of $V$ is far more pronounced. Fixing $U$ at 1.5 eV and varying $V$ from 0.5 to 1.5 eV, we find that the low-energy band structure is highly sensitive to $V$. As $V$ increases, the M-shaped nature of the SP1 band is completely disrupted, and the relative position of the SP1 and SP2 bands changes significantly.
The band dispersion of the ISD phase across the entire Brillouin zone (calculated using the cRPA-derived values: $U$ = 1.35 eV and $V$ = 0.17 eV), is provided in Fig. S6 Sup , along with that of the pristine phase for comparison.
The positions of the van Hove singularities are essential for identifying electronic instabilities Li et al. (2021); Ferrari et al. (2022); Xie et al. (2022). In the pristine phase, one of the prominent van Hove singularities appears near the M-point. This feature is found to be split in the ISD phase, as demonstrated in Fig. S6 Sup . The split bands are associated with the M-shaped band shown in Fig. 4(a). The dominant orbital contributions to the M-shaped band arise from the $3d_{x^{2}-y^{2}}$, $3d_{xy}$, and $3d_{z^{2}}$ orbitals, which give rise to pronounced peaks near -0.2 eV in the density of states (see Fig. 5). Note that the van Hove singularity peaks are substantially suppressed in the ISD phase relative to the pristine phase, due to the band splitting near the M-point.
Now We account for structural contributions corresponding to different $U$ and $V$ values, as discussed in Sec. III.2, which were neglected in the previous calculations. Figure S5 (in the Supplementary Material Sup ) shows the band-unfolding results for the optimized $2\times 2$ ISD structures with corresponding $U$ and $V$ values. For higher $V$ values, structural optimization induces large atomic displacements of vanadium atoms, resulting in significant energy band splitting. Notably, in the absence of $V$, the M-shaped nature of the SP1 band is almost entirely suppressed as $U$ increases from 0 to 2 eV. This suggests that the unfolded band dispersion closely resembles that of the pristine structure (shown in Fig. S4 Sup ), reinforcing our earlier analysis that CDW deformations diminish and the system transitions towards a stable pristine structure as $U$ increases. It is worth noting that the primary effect of the structural contributions is the modification of the band dispersion arising from the larger atomic distortions at finite $U$ and $V$ values.
III.5 Electronic structures: density of states
@@
To gain deeper insight into the electronic structure of KV3Sb5, we calculated the density of states (DOS) based on the local axis, as shown in Fig. 5(d). In the $2\times 2$ ISD phase, vanadium atoms occupy two inequivalent sites in the kagome lattice, denoted as V1 and V2 (Fig. 1). The orbital-resolved DOS for V1 and V2 in the $2\times 2$ ISD phase is shown in Fig. 5. To examine the effects of $U$ and $V$ solely, we used the optimized ISD structure obtained at $U=0$ and $V=0$. The DOS for V1 and V2 exhibit similar behavior, with V1 contributing slightly more than V2 near the Fermi level, which remains unchanged for all values of $U$ and $V$. Furthermore, in the absence of $V$, increasing $U$ enhances the DOS at the Fermi level, whereas incorporating $V$ (without $U$) leads to the opposite trend. Specifically, $V$ suppresses the DOS at the Fermi level, except for the $d_{xz/yz}$ orbitals, which contribute to the $\Lambda$-shaped (SP2) band, as shown in the projected band dispersion in Fig. S7 Sup . This interplay between $U$ and $V$ in shaping the PDOS at the Fermi level is illustrated in Fig. S8 of the Supplementary Material Sup . It is well established that CDW phases disrupt van Hove singularities, leading to a reduction in the DOS at the Fermi level Tan et al. (2021). These finding align with our earlier results in Sec. III.2, where we observed that increasing $U$ destabilizes the CDW phase, making the pristine structure more favorable, while higher values of $V$ help sustain the CDW phase.
We further calculated the $3d$ occupation numbers for V1 and V2 across different values of $U$ and $V$. These were determined by integrating the DOS up to the Fermi energy in the valence band. Figure 5(c) shows the occupation numbers for V1 and V2 as a function of $U$ for different $V$ values in the $2\times 2$ ISD phase. For $U=0$ and $V=0$, the occupation number of V1 (V2) is nearly 4.102 (4.093), decreasing to 4.055 (4.056) at $U=3$ eV and $V=0$. Our results confirm that for any given $V$, the occupation number decreases with increasing $U$, whereas for a fixed $U$, increasing $V$ leads to the opposite trend. The orbital-decomposed occupation numbers, shown in Fig. S9 Sup , further illustrate this contrasting behavior of $U$ and $V$.
@@ IV Conclusions
In summary, we systematically investigated the roles of on-site ($U$) and inter-site ($V$) Coulomb interactions in shaping the charge-density-wave (CDW) phases of the kagome metal KV3Sb5. From our DFT+$U$+$V$ calculations, we constructed the $U$-$V$ phase diagram, revealing the contrasting roles of $U$ and $V$ in forming CDW phases. We found that $V$ plays a decisive role in stabilizing the inverse Star of David (ISD) CDW phase, consistent with experimental observations. The interplay between $U$ and $V$ influences not only the relative stability and extent of lattice distortions but also modulates the electronic instabilities associated with van Hove singularities near the Fermi level. By examining unfolded band structures, we clearly demonstrated how variations in $U$ and $V$ modify the band shapes and dispersions. Our cRPA estimates provided realistic values of $U$ and $V$, enabling us to identify the position of KV3Sb5 in the $U$-$V$ phase space. Through comparison with experimental ARPES data, we demonstrated that our chosen interaction parameters accurately reproduce the band structure of the system.
Our findings underscore the importance of both local and nonlocal electronic correlations in kagome systems and provide a unified framework for understanding the emergence of exotic phases—including unconventional superconductivity and topological phenomena—in kagome metals. This work offers valuable insights for future studies of correlated states in kagome-lattice materials.
@@ Acknowledgements
BK acknowledges support from NRF (NRF2021R1A4A1031920, RS-2021-NR061400, and RS2022-NR068223) and KISTI Supercomputing Center (Project No. KSC-2023-CRE-0413). CJK was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean Government (MSIT) (Grant No. 2022R1C1C1008200).

### v2

@@ III.5 Electronic structures: band structure
Figure 4: Calculated unfolded band dispersion for $2\times 2$ ISD phase along K-M-K path for different values of $U$ and $V$ (in eV): (a) cRPA values ($U=1.35,V=0.17$), (b) $U=0$, $V=0$, (c) $U=1$, $V=0$, (d) $U=1.5$, $V=0$, (f) $U=1.5$, $V=0.5$, (g) $U=1.5$, $V=1$, and (h) $U=1.5$, $V=1.5$. (e) ARPES intensity measured at $T$ = 20 K along the K-M-K cut at $h\nu=114$ eV, as adapted from Kato et al. [37]. The black solid lines in (a) are a guide for the eye to understand both the SP1 and SP2 bands. A Fermi level offset of 200 meV is introduced in the DFT+UU+$V$ bands to enable direct comparison with the experimental spectra.
We now investigate the electronic structure of KV3S5, focusing on its dependence on the Hubbard parameters $U$ and $V$. Given the weak interlayer interactions, we consider the $2\times 2$ ISD phase, as shown in Fig. 1. To isolate the effects of $U$ and $V$ from structural contributions, we use the optimized ISD structure obtained at $U=0$ and $V=0$ as the baseline for calculations with various $U$ and $V$ values. The phase shifts between adjacent kagome layers have minimal impact on the band dispersion. Thus, our approach is sufficient for investigating the electronic structure of KV3S5. Structural contributions, such as variations in vanadium atom distortion for different values of $U$ and $V$, will be discussed at the end of this subsection. Figure 4 presents the band dispersion of the $2\times 2$ ISD phase for various values of $U$ and $V$, unfolded onto the $1\times 1$ Brillouin zone. The $k$-path K-M-K is selected to facilitate direct comparison with existing angle-resolved photoemission spectroscopy (ARPES) measurements, as shown in Fig. 4(e). The band dispersion of the ISD phase successfully captures two dominant features near the Fermi level: (i) an M-shaped band and (ii) a $\Lambda$-shaped band, as highlighted by the guidelines in Fig. 4(a). These features are consistent with the experimental ARPES measurement by Kato et al. [37], where the former is denoted as the SP1 band and the latter as SP2. Near the M-point, the $\Lambda$-shaped (SP2) band penetrates the M-shaped (SP1) band and reaches a higher binding energy in the ISD phase. In contrast, no such penetration occurs in the SD phase, where the $\Lambda$-shaped (SP2) band remains at a higher energy position. We found that our on-site and inter-site Hubbard parameters, $U$ = 1.35 eV and $V$ = 0.17 eV, obtained from the cRPA calculations, successfully reproduce the key features of the experimental ARPES described above (Fig. 4 (a)). Both the relative energy positions of the SP1 and SP2 bands and their morphology closely match the experimental spectra in Fig. 4(e). This agreement contrasts with the results obtained using other $U$ and $V$ values, as shown in the Fig. 4. While DFT calculations without $U$ and $V$ (Fig. 4(b)) also show reasonable agreement with ARPES data, the cRPA-derived finite $U$ and $V$ values play a crucial role in shaping the band dispersion of the KV3Sb5 system, as demonstrated in Fig. S5 of the Supplementary Material [65]. Our systematic investigation reveals the distinct roles of $U$ and $V$ in shaping the low-energy band structures. Variations in $U$ primarily affect the M-shaped (SP1) band, which predominantly has $3d_{x^{2}-y^{2}}$, $3d_{xy}$, and $3d_{z^{2}}$ orbital characters, while the $\Lambda$-shaped (SP2) band, mainly associated with $3d_{xz/yz}$ orbital characters, remains nearly rigid (see Fig. S7 in the Supplementary Material [65] for detailed orbital characters). Note that the orbital characters of both SP1 and SP2 bands are consistent with the earlier findings by Kato et al. [37]. Increasing $U$ progressively shifts the SP1 band downward, lowering its energy at the $K$-point, whereas $U$ has minimal impact on the position of the SP2 band. In contrast, the effect of $V$ is far more pronounced. Fixing $U$ at 1.5 eV and varying $V$ from 0.5 to 1.5 eV, we find that the low-energy band structure is highly sensitive to $V$. As $V$ increases, the M-shaped nature of the SP1 band is completely disrupted, and the relative position of the SP1 and SP2 bands changes significantly. The band dispersion of the ISD phase across the entire Brillouin zone (calculated using the cRPA-derived values: $U$ = 1.35 eV and $V$ = 0.17 eV), is provided in Fig. S6 [65], along with that of the pristine phase for comparison. The positions of the van Hove singularities are essential for identifying electronic instabilities [25, 49, 42]. In the pristine phase, one of the prominent van Hove singularities appears near the M-point. This feature is found to be split in the ISD phase, as demonstrated in Fig. S6 [65]. The split bands are associated with the M-shaped band shown in Fig. 4(a). The dominant orbital contributions to the M-shaped band arise from the $3d_{x^{2}-y^{2}}$, $3d_{xy}$, and $3d_{z^{2}}$ orbitals, which give rise to pronounced peaks near -0.2 eV in the density of states (see Fig. 5). Note that the van Hove singularity peaks are substantially suppressed in the ISD phase relative to the pristine phase, due to the band splitting near the M-point. Now we account for structural contributions corresponding to different $U$ and $V$ values, as discussed in Sec. III.3, which were neglected in the previous calculations. Figure S5 (in the Supplementary Material [65]) shows the band-unfolding results for the optimized $2\times 2$ ISD structures with corresponding $U$ and $V$ values. For higher $V$ values, structural optimization induces large atomic displacements of vanadium atoms, resulting in significant energy band splitting. Notably, in the absence of $V$, the M-shaped nature of the SP1 band is almost entirely suppressed as $U$ increases from 0 to 2 eV. This suggests that the unfolded band dispersion closely resembles that of the pristine structure (shown in Fig. S4 [65]), reinforcing our earlier analysis that CDW deformations diminish and the system transitions towards a stable pristine structure as $U$ increases. It is worth noting that the primary effect of the structural contributions is the modification of the band dispersion arising from the larger atomic distortions at finite $U$ and $V$ values.
III.6 Electronic structures: density of states
@@
To gain deeper insight into the electronic structure of KV3Sb5, we calculated the density of states (DOS) based on the local axis, as shown in Fig. 5(d). In the $2\times 2$ ISD phase, vanadium atoms occupy two inequivalent sites in the kagome lattice, denoted as V1 and V2 (Fig. 1). The orbital-resolved DOS for V1 and V2 in the $2\times 2$ ISD phase is shown in Fig. 5. To examine the effects of $U$ and $V$ solely, we used the optimized ISD structure obtained at $U=0$ and $V=0$. The DOS for V1 and V2 exhibit similar behavior, with V1 contributing slightly more than V2 near the Fermi level, which remains unchanged for all values of $U$ and $V$. Furthermore, in the absence of $V$, increasing $U$ enhances the DOS at the Fermi level, whereas incorporating $V$ (without $U$) leads to the opposite trend. Specifically, $V$ suppresses the DOS at the Fermi level, except for the $d_{xz/yz}$ orbitals, which contribute to the $\Lambda$-shaped (SP2) band, as shown in the projected band dispersion in Fig. S7 [65]. This interplay between $U$ and $V$ in shaping the PDOS at the Fermi level is illustrated in Fig. S8 of the Supplementary Material [65]. It is well established that CDW phases disrupt van Hove singularities, leading to a reduction in the DOS at the Fermi level [36]. These finding align with our earlier results in Sec. III.3, where we observed that increasing $U$ destabilizes the CDW phase, making the pristine structure more favorable, while higher values of $V$ help sustain the CDW phase. We further calculated the $3d$ occupation numbers for V1 and V2 across different values of $U$ and $V$. The occupation numbers were determined by integrating the partial DOS of the $d$-orbitals up to the Fermi energy. Figure 5(c) shows the occupation numbers for V1 and V2 as a function of $U$ for different $V$ values in the $2\times 2$ ISD phase. For $U=0$ and $V=0$, the occupation number of V1 (V2) is nearly 4.102 (4.093), decreasing to 4.055 (4.056) at $U=3$ eV and $V=0$. Our results confirm that for any given $V$, the occupation number decreases with increasing $U$, whereas for a fixed $U$, increasing $V$ leads to the opposite trend. The orbital-decomposed occupation numbers, shown in Fig. S9 [65], further illustrate this contrasting behavior of $U$ and $V$.
@@ IV Conclusions
In summary, we systematically investigated the roles of on-site ($U$) and inter-site ($V$) Coulomb interactions in shaping the charge-density-wave (CDW) phases of the kagome metal KV3Sb5. From our DFT+UU+$V$ calculations, we constructed the $U$-$V$ phase diagram, revealing the contrasting roles of $U$ and $V$ in forming CDW phases. We found that $V$ plays a decisive role in stabilizing the inverse Star of David (ISD) CDW phase, consistent with experimental observations. The interplay between $U$ and $V$ influences not only the relative stability and extent of lattice distortions but also modulates the electronic instabilities associated with van Hove singularities near the Fermi level. By examining unfolded band structures, we clearly demonstrated how variations in $U$ and $V$ modify the band shapes and dispersions. Our cRPA estimates provided realistic values of $U$ and $V$, enabling us to identify the position of KV3Sb5 in the $U$-$V$ phase space. Through comparison with experimental ARPES data, we demonstrated that our chosen interaction parameters accurately reproduce the band structure of the system. Our findings underscore the importance of both local and nonlocal Coulomb interactions in kagome systems and provide a unified framework for understanding the emergence of exotic phases—including unconventional superconductivity and topological phenomena—in kagome metals. This work offers valuable insights for future studies of correlated states in kagome-lattice materials.
@@ Acknowledgements
BK acknowledges support from NRF (NRF2021R1A4A1031920, RS-2021-NR061400, and RS2022-NR068223) and KISTI Supercomputing Center (Project No. KSC-2023-CRE-0413). CJK was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean Government (MSIT) (Grant No. 2022R1C1C1008200). The authors acknowledge the support from the Advanced Study Group program from PCS-IBS and the hospitality at APCTP, where part of this work was done.